If C dollars are deposited in an account paying r percent annual interest, approximate the amount in the account after x years if C= $150, r= 2.2%,and x= 19 years.
A= $....... round to the nearest cent

Answer by mathstutor458(57)

C= $150, r= 2.2%,and x= 19 years.
Simple interest = C*r*x/100 = 150*2.2*19/100=$62.70
Therefore, Amount=C+interest = 150+ 62.7=$212.70